OpenStudy (anonymous):

Which expression is equivalent to 6 – (–8)?

OpenStudy (anonymous):

@Smileyside

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

–6 + 8 –6 + –8 6 + –8 6 + 8

OpenStudy (butterflydreamer):

think of it as 6 + ( -1 * -8) So when we multiply -1 * -8 , we get?

OpenStudy (anonymous):

8 using my calculator

OpenStudy (butterflydreamer):

correct so 6 - ( -8) = 6 + ( -1 * -8) = 6 + (8) = 6 + 8

OpenStudy (anonymous):

14

OpenStudy (butterflydreamer):

that's true xD BUT look at your options that you can choose from.. .

OpenStudy (anonymous):

i know what it is now..

OpenStudy (butterflydreamer):

greaaat ^_^!